OpenStudy (anonymous):

r: (h, k) = (0, 0), r = 4 Find the equation of the circle in standard form for the given center (h, k) and radius

OpenStudy (anonymous):

Do you know what the standard form of a circle looks like?

OpenStudy (anonymous):

yes please

OpenStudy (princeharryyy):

(x-h)^2 + (y-k)^2 = r^2

OpenStudy (anonymous):

\[(x−h)^2+(y−k)^2=r^2\]

OpenStudy (princeharryyy):

replace h,k,r and solve

OpenStudy (princeharryyy):

@Hannah777

OpenStudy (princeharryyy):

so, here (x-0)^2 + (y-0)^2 + 4^2 => x^2 + y^2 = 16 is the required equation @Hannah777